solve -8(x-3)+12x=15-(2x+3)
-8*x+24+12*x=15-2*x-3
6x+12

Found 2 solutions by jst_kidding, asad:
Answer by jst_kidding(6)

-8(x-3)+12x=15-(2x+3)
-8x+24+12x=15-2x-3
4x+24=12-2x
6x=-12
x=-2

Answer by asad(2)